OpenStudy (anonymous):

assume f (x)= -2x+8 and g (x)=5x. What is the value of (f o g) (3)?

OpenStudy (mertsj):

f o g=-2(5x)+8=-10x+8 (f o g) (3)=-10(3)+8=-22